# ViSurf: Visual Supervised-and-Reinforcement Fine-Tuning for Large Vision-and-Language Models

### v2

@@ 3.3 Combining SFT and RLVR
Gradient Analysis of ViSurf. The gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We omit clip operation for simplicity:
$\nabla_{\theta}\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (9)
$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)$
$\qquad+{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Bigg)\Bigg]_{\,\theta\approx\theta_{\mathrm{old}}}.$
Relation to SFT and RLVR. We analyze the gradients of ViSurf, SFT, and RLVR to explain their connections. To better illustrate the structure of the gradient, we reformulate Equation 9 as following:
@@
${\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\Biggl[\frac{1}{G+1}\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\text{RLVR Term}}}}$
@@
${\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\Biggl[\frac{1}{G+1}{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\text{SFT Term}}}}.$
@@
The RLVR term in Equation 10 is structurally identical to the standard RLVR gradient in Equation 5, differing only in its scaling coefficient ($\frac{1}{G+1}\hat{A}_{j}$ vs. $\frac{1}{G}\hat{A}_{j}$). Similarly, the SFT term in Equation 10 resembles the SFT gradient from Equation 4, with two key distinctions: (i) the coefficient is weighted by $\frac{1}{G+1}\hat{A}_{y}$ instead of 1, and (ii) the use of the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$. This approximation implies that the ground-truth label $y$ should align with the model’s internal generative preference to be effective. Crucially, Equation 9 integrates both the external guidance from SFT and the internal guidance from RLVR.
3.4 Reward Control for Ground-Truth Label
The advantage $\hat{A}y$ for the ground-truth label $y$ is always positive till now, as the label is correct and receives higher reward, which can lead to reward hacking. The setup is sub-optimal for two reasons: (i) the ground-truth lacks a reasoning trace, and (ii) it suppresses the relative advantage $\hat{A}_{j}$ even if rollouts have already generated correct trace and answer. Furthermore, as analysis above, we must ensure the ground-truth aligns with the self-rollouts to satisfy the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$. To address these issues, we propose three reward control strategies.
Aligning Ground-truth Labels with Rollouts Preference. To ensure compatibility between the ground-truth data and the self-rollouts generated by $\pi_{\theta}$, we reformat the ground-truth annotations to match the model’s preferred output style. For instance, we adjust the whitespace in JSON-like structures from {‘bbox’:[x1,y1,x2,y2]} to {‘bbox’:x[x1,xy1,xx2,xy2]}(i.e., adding space after the punctuation), as these variants yield different tokenizations. This alignment minimizes the distribution shift between $\pi_{\theta}$ and $\pi_{\theta_{old}}$, satisfying the underlying assumption $\pi_{\theta}\approx\pi_{\theta_{old}}$.
Eliminating Thinking Reward for Ground-truth Labels. Since the ground-truth labels lack annotated reasoning path, we assign a reasoning format score of zero to them. This operation ensures the model learns reasoning trace directly from its self-rollouts without being biased by missing external reasoning annotations.
Smoothing the Reward for Ground-truth Labels. Prior to advantage estimation, we compare the maximum reward among generated rollouts, $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$, against the ground-truth reward $\text{r}(y)$. If $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}\geq\text{r}(y)$, it indicates the policy model $\pi_{\theta}$ has already produced a high-quality output without external guidance. In this case, we set $\text{r}(y)=\text{mean}\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$. This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), eliminating the external supervision signal when it is unnecessary.
3.5 Optimization Analysis During Training
Upper Bound Analysis. As mentioned above, our ViSurf is particularly beneficial when old policy model $\pi_{\theta_{old}}$ cannot generate correct rollouts. When the old policy model $\pi_{\theta_{old}}$ already achieves desirable rollouts, the SFT Term in Equation 10 near to zero, thus the upper bound of ViSurf is the RLVR. However, when the policy model cannot generate desirable rollouts, the upper bound is better than using either SFT or RLVR alone.
4 Experiments
@@
[…]
@@ 4.1 Experimental Settings
We verify ViSurf on benchmarks across several domains.
Non-Object Segmentation. The gRefCOCO [liu2023gres] includes queries that do not contain corresponding objects. The evaluation metrics are gIoU and N-Acc. We use Multi-objects-7K plus with 200 non-object data for training.
Reasoning Segmentation. The ReasonSeg [lai2024lisa] includes test samples that need reasoning for correct segmentation. It has 200 validation images and 779 test images. The evaluation metric is gIoU. We use Multi-objects-7K proposed in VisionReasoner [liu2025visionreasoner] for training.
GUI Grounding. The OmniACT [kapoor2024omniact] is a GUI grounding task for Desktop and Web. We derive 6,101 samples in training split and verify on the test split. The accuracy is calculated as whether the predict point correctly locates inside the interest region.
Anomaly Detection. The RealIAD [wang2024real] includes real-world, multi-view industrial anomaly. We derive 3,292 training samples and 2,736 test samples, ensuring the two sets are disjoint. We calculated the ROC_AUC.
Medical Image: Skin. The task one of ISIC2018 [codella2019skin, isic2018] is lesion segmentation. It includes 2,594 training samples and 1,000 test samples. We measure the bbox_acc metric, which computes the ratio of predicted bounding boxes whose IoU with the ground truth exceeds 0.5.
MathVista. The MathVista-testmini [lu2023mathvista] includes 1,000 diverse mathematical and visual tasks. We gather around 10k training data from WeMath [qiao2024we], MathVision [wang2024measuring], Polymath [gupta2024polymath], SceMQA [liang2024scemqa], Geometry3K [lu2021inter].
Implemention Details. We instantialize ViSurf algorithm with Qwen2.5VL-7B [bai2025qwen25vl] and adopt SAM2 [ravi2024sam2] if needed. We employ a constant learning rate of 1e-6 for all methods, with a batch size of 32 for SFT and 16 for RLVR and ViSurf. We employ same training steps for fair comparison. For MathVista, the reward function consists of format and accuracy rewards. For other tasks, we adopt the rewards from VisionReasoner [liu2025visionreasoner], which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
@@ 4.2 Comparison of Different Training Paradigms
Main Results. A comparative analysis of post-training paradigms across various domains is presented in Table 1. Empirical results indicate that ViSurf consistently outperforms existing post-training paradigms across various domains, with an average relative gain of 38.6% over the baseline model. This advantage is particularly significant in domains where the baseline model demonstrates lower competency, such as Non-Object and Anomaly, suggesting the method’s efficacy in addressing domain exceeding model’s knowledge base. However, in domains where the baseline model is already highly proficient, the incremental gains are relatively marginal. We also observe that SFT leads to performance degradation in OmniACT, which may be attributable to potential ‘test data contamination’ during baseline model’s pre-training stage. In comparison, both RLVR and ViSurf preserve the baseline model’s performance. Furthermore, in the case of RealIAD and non-object detection in gRefCOCO, the pure RLVR approach underperforms relative to the original model. We attribute this to the fact that self-generated rollouts frequently produce incorrect answers, thereby hindering model optimization.
Catastrophic Forgetting. We evaluate the performance of ChartQA [masry2022chartqa] and DocVQA [mathew2021docvqa]. As illustrated in Table 2, VQA performance exhibits notable variation across different training paradigms. Both RLVR and ViSurf demonstrate robustness against catastrophic forgetting. In contrast, SFT and SFT → RLVR suffer from performance degradation, which is attributable to catastrophic forgetting.
@@
[…]
ViSurf on Other Models. We apply ViSurf to the Qwen2VL-7B [wang2024qwen2vl]. As shown in Table 3, our method consistently outperforms its counterparts. Furthermore, the pure RLVR approach yields the weakest performance on both datasets and even underperforms the baseline on RealIAD, which indicates that external supervision is critical.
@@ 4.3 Ablation of Reward Control
Eliminating Thinking Reward for Ground-truth Labels. The results indicate that the reasoning strategy is critical for tasks requiring complex inference, such as those in ReasonSeg and MathVista, as it encourages the model to generate a reasoning process prior to delivering the final answer. Conversely, for the gRefCOCO dataset, where queries are typically limited to simple class types (e.g., ”human”) and basic references (e.g., ”woman on the right”), omitting the reasoning step yields superior performance. This suggests that the necessity of explicit reasoning is contingent upon the complexity of the underlying task.
@@ 4.4 In-depth Analysis
Reduce the Burden of Prompt Design. The RLVR paradigm relies heavily on explicit instructions to guide the model toward generating rollouts in a specific format, such as output with format ’point_2d’: [2, 3]. In contrast, ViSurf incorporates external guidance with desired format, thereby reducing the dependency on manual prompt engineering. Table 5 compares performance with and without detailed prompts (details are provided in Appendix D of the supplementary materials), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
@@
[…]
Training Cost. We conducted a comparative analysis of the per-step training time for different fine-tuning paradigms. Each method was implemented using well-established frameworks: DeepSpeed [deepspeed] and TRL [trl] for SFT, and VeRL [verl] for RLVR and ViSurf. The results indicate that while RLVR and ViSurf offer greater memory efficiency, they incur a higher computational cost per step, attributable to the overhead of generating rollouts.

### v3

@@ 3.3 Combining SFT and RLVR
Gradient Analysis of ViSurf. In our practical implementation, we adopt a single-sampling, single-updating paradigm, which effectively bypasses the need for clipping operations. Therefore, the gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick:
$\nabla_{\theta}\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (9)
$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)$
$\qquad+{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Bigg)\Bigg]_{\,\theta\approx\theta_{\mathrm{old}}}.$
3.4 Reward Control for Ground-Truth Label
The advantage $\hat{A}y$ for the ground-truth label $y$ is always positive till now, as correct labels inherently receive higher rewards. However, this static setup is often sub-optimal; it can lead to reward hacking and suppresses the relative advantage $\hat{A}_{j}$ of self-generated rollouts, even when the policy has already produced correct trace and answers. Furthermore, maintaining the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ requires ensuring that ground-truth data distributions remain aligned with self-generated rollouts. To mitigate these issues, we propose three reward control strategies:
Aligning Ground-truth Labels with Rollouts Preference. To ensure compatibility between the ground-truth annotations and the outputs generated by $\pi_{\theta}$, we reformat the ground-truth data to match the model’s preferred stylistic patterns. For instance, we adjust the whitespace in JSON-like structures from {‘bbox’:[x1,y1,x2,y2]} to {‘bbox’:x[x1,xy1,xx2,xy2]} (i.e., adding space after the punctuation), because minor variations in punctuation can lead to significantly different tokenizations. This alignment minimizes the distribution shift between $\pi_{\theta}$ and $\pi_{\theta_{old}}$, thereby upholding the core assumptions $\pi_{\theta}\approx\pi_{\theta_{old}}$.
Eliminating Thinking Reward for Ground-truth Labels. Since the ground-truth labels lack annotated reasoning path, we assign a reasoning format score of zero to them when reasoning is needed. This ensures that the model learns to generate its own reasoning traces through self-rollouts without being biased or penalized by the absence of external reasoning annotations in the ground-truth data.
Smoothing the Reward for Ground-truth Labels. Prior to advantage estimation, we compare the maximum reward among generated rollouts, $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$, against the ground-truth reward $\text{r}(y)$. If $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}\geq\text{r}(y)$, it indicates the policy model $\pi_{\theta}$ has already produced a high-quality output without external guidance. In this case, we set $\text{r}(y)=\text{mean}\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$. This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), eliminating the external supervision signal.
3.5 Optimization Analysis During Training
To better analysis the optimization step, we reformulate the gradient Equation 9 as following:
@@ 3.5 Optimization Analysis During Training
${\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\Biggl[\frac{1}{G+1}\sum_{j=1}^{G}{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\text{RLVR Term}}}}$
@@
${\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\Biggl[\frac{1}{G+1}{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\text{SFT Term}}}}.$
@@
The RLVR term in Equation 10 is structurally identical to the standard RLVR gradient in Equation 5, differing only in its scaling coefficient ($\frac{1}{G+1}\hat{A}_{j}$ vs. $\frac{1}{G}\hat{A}_{j}$). Similarly, the SFT term in Equation 10 resembles the SFT gradient from Equation 4, with two key distinctions: (i) the coefficient is weighted by $\frac{1}{G+1}\hat{A}_{y}$ instead of 1, and (ii) the use of the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$. Crucially, Equation 9 integrates both the external guidance from SFT and the internal guidance from RLVR.
@@
ViSurf 66.6 57.1 66.5 65.0 65.6 69.3 94.7 71.6 69.6
Upper Bound Analysis. As mentioned above, our ViSurf is particularly beneficial when old policy model $\pi_{\theta_{old}}$ cannot generate correct rollouts. When the old policy model $\pi_{\theta_{old}}$ already achieves desirable rollouts, the SFT Term in Equation 10 equals to zero, thus the upper bound of ViSurf is the RLVR. However, when the policy model cannot generate desirable rollouts, the upper bound is better than using either SFT or RLVR alone.
4 Experiments
@@ 4.1 Experimental Settings
We verify ViSurf on benchmarks across several domains, including Non-Object Segmentation (e.g., gRefCOCO (Liu et al., 2023a)), Reasoning Segmentation (e.g., ReasonSeg (Lai et al., 2024)), GUI Grounding (e.g., OmniACT (Kapoor et al., 2024)), Industrial Anomaly Detection (e.g., RealIAD (Wang et al., 2024a)), Medical Imaging (e.g., ISIC2018 (Codella et al., 2019)), and Mathematical Reasoning (e.g., MathVista (Lu et al., 2023)). Details of data split and evaluation protocols are provided in the Appendix E.
Implemention Details. We instantialize ViSurf algorithm with Qwen2.5VL-7B (Bai et al., 2025) and adopt SAM2 (Ravi et al., 2024) if needed. We employ a constant learning rate of 1e-6 for all methods, with a batch size of 32 for SFT and 16 for RLVR and ViSurf. We employ same training steps for fair comparison. For MathVista, the reward function consists of format and accuracy rewards. For other tasks, we adopt the rewards from VisionReasoner (Liu et al., 2025c), which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
@@ 4.2 Comparison of Different Training Paradigms
Main Results. A comparative analysis of various post-training paradigms is summarized in Table 1. Empirical evaluations demonstrate that ViSurf consistently surpasses existing methodologies across all evaluated domains, achieving a substantial average relative improvement of 38.6% over the baseline. This performance gain is most pronounced in challenging domains such as Non-Object and Anomaly, where the baseline initially struggles. This suggests that ViSurf is particularly effective at expanding a model’s capabilities in areas that exceed its inherent knowledge base. Conversely, in domains where the baseline already exhibits high proficiency, we observe more marginal incremental improvements. Our analysis also reveals that standard Supervised Fine-Tuning (SFT) leads to performance degradation on OmniACT, a phenomenon likely stemming from overfitting during the baseline’s original pre-training phase. In contrast, both RLVR and ViSurf successfully preserve the baseline’s foundational performance. Notably, in RealIAD and gRefCOCO (non-object detection), the pure RLVR approach actually underperforms the original model. We attribute this to the high frequency of incorrect self-generated rollouts, which introduce noise and hinder effective model optimizationm, a bottleneck that ViSurf effectively mitigates.
Catastrophic Forgetting. We evaluate the performance of ChartQA (Masry et al., 2022) and DocVQA (Mathew et al., 2021) after fine-tuning without VQA data. As illustrated in Table 2, VQA performance exhibits notable variation across different training paradigms. Both RLVR and ViSurf demonstrate robustness against catastrophic forgetting. In contrast, SFT and SFT → RLVR suffer from performance degradation, which is attributable to catastrophic forgetting.
@@
   ViSurf    76.0    95.4
ViSurf on Other Models. We apply ViSurf to the Qwen2VL-7B (Wang et al., 2024c). As shown in Table 3, our method consistently outperforms its counterparts.
@@ 4.3 Ablation of Reward Control
Eliminating Thinking Reward for Ground-truth Labels. The results indicate that the reasoning strategy is critical for tasks requiring complex inference, such as those in ReasonSeg and MathVista, as it encourages the model to generate a reasoning process prior to delivering the final answer. Conversely, for the gRefCOCO dataset, where queries are typically limited to simple class types (e.g., ”human”) and basic references (e.g., ”woman on the right”), omitting the reasoning trace yields superior performance. This suggests that the necessity of explicit reasoning is contingent upon the complexity of the underlying task.
@@ 4.4 In-depth Analysis
Reduce the Burden of Prompt Design. The RLVR paradigm relies heavily on explicit instructions to guide the model toward generating rollouts in a specific format, such as output with format ’point_2d’: [2, 3]. In contrast, ViSurf incorporates external guidance with desired format, thereby reducing the dependency on manual prompt engineering. Table 5 compares performance with and without detailed prompts (prompts difference are shown in the supplementary materials), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
@@
ViSurf 81.8 22.9
Training Cost. We conducted a comparative analysis of the per-step training for different fine-tuning paradigms. Each method is implemented using well-established frameworks: DeepSpeed (Microsoft and DeepSpeed Team, 2020) and TRL (Face, 2024) for SFT, and VeRL (Seed, 2024) for RLVR and ViSurf. The results indicate that while RLVR and ViSurf offer greater memory efficiency, they increase computational time, attributable to the overhead of generating rollouts.